The following information is available.
H0: μ ≤ 25
H1: μ > 25
The sample mean is 26 for a sample of 46. The population standard deviation is 2. Use the .01 significance level. What is the p-value?

Answer by ewatrrr(24785)

H0: μ ≤ 25
H1: μ > 25
Sample m = 26, SD = 2, n = 26
t(26) = (26-25)/[2/sqrt(26)] = 1/.3922 = 2.55
p-value = P(t> 2.55 when df = 25) = 1 - tcdf(2.55,100,25) = 1 - 0.9914 = .0086
At the .01 significance level, .0086 < .01, Reject Ho.